Celsius to Fahrenheit. Water freezes at 0 degrees C or 32 degrees F and boils at 100 degrees C or 212 degrees F. There is a linear equation that expresses the number of degrees Fahrenheit (F) in terms of the number of degrees Celsius (C). Find the equation and find the Fahrenheit temperature when the Celsius temperature is 45 degrees.
Answer by rfer(10417)

F=9/5*C+32
F=9/5*45+32
F=81+32
F=113 degrees F